OpenStudy (anonymous):

use the formula method to solve the equation 21x^2+17x-250=0

OpenStudy (anonymous):

-(17)+-sqrt (17)^2 -4(21)(-250)/2(21) then simplify and answer

OpenStudy (anonymous):

21x^2 + 17x - 250 = 0 Using the quadratic formula we get: x = (- 17 ± √ [17^2 + 4(21)(250)])/42 => x = ( -17 ± √ 21289)/42 i.e. x = 3.07 so, total time is 4x + 2 = 14.3 seconds